# Measurement-device agnostic quantum tomography

### v2

@@ Abstract
The tomography consists of projective measurements. First, the analyzed state undergoes unitary evolution provided by a half- and quarter-wave plate and is subsequently projected using a horizontally oriented linear polarizer. The action of the waveplate could be seen as the rotation of the Bloch sphere around axis $(\sin(2\alpha),0,\cos(2\alpha))$ with rotation angle $\Gamma$, where $\alpha$ determines the angular position of the waveplate and $\Gamma$ its retardance. The retardance usually slightly deviates from its nominal value. These deviations manifest as under- or over-rotation of the Bloch sphere. We parameterize the measurement operator by precisely controlled waveplate angular positions $x$, $y$, and half- and quarter-wave plate retardance deviations $\delta$ and $\epsilon$, respectively. These deviations are initially unknown to us, and we will use the presented method to estimate their values.
@@
To demonstrate the tolerance of the method to a unitary operation between preparation and tomographic measurement, we add a quarter-wave plate with its fast axis oriented at -12.5 deg relative to the horizontal direction. We use eight probe states and perform the Pauli tomography for each. Initially, we assume zero deviations, $\delta=\epsilon=0$, and observe the purity modulation $\Delta P=0.06$. Then, we vary parameters $\delta,\epsilon$ and update the measurement operators employed in the reconstruction to minimize the purity modulation. The optimal $\delta$ and $\epsilon$ are close to the true values because the reconstruction artifact is significantly reduced to $\Delta P=0.01$. We found the optimal parameters $\delta=5.5$ deg, $\epsilon=-1.5$ deg.
@@
We then reverse the process to determine retardance deviations $\tilde{\delta},\tilde{\epsilon}$ of the waveplates used in state preparation. We prepare the six eigenstates of Pauli operators, project each onto eight states corresponding to the original probe state set, and treat this data as before. We minimize the apparent purity modulation to find the optimal parameters $\tilde{\delta}=4.5$ deg and $\tilde{\epsilon}=-3.6$ deg, reducing the purity modulation from 0.07 to 0.01.
We verified the accuracy of the determined parameters using the following steps. We removed the central quarter-wave plate and then prepared an ensemble of test states, depicted as empty circles in Figure 2. As a reference, we first assumed perfect waveplates. The minimum fidelity of the reconstructed state to the target states was $F_{\min}=0.967$, and the average fidelity was $\bar{F}=0.992$. It indicates imprecise preparation. The purity modulation, $\bar{P}=0.990$ and $P_{\min}=0.938$, indicates reconstruction artifacts. When we used our knowledge of $\delta$ and $\epsilon$ in the reconstruction, the average purity increased to $\bar{P}=0.995$ and the minimum purity increased to $0.982$. The information gained in the calibration clearly mitigated the reconstruction artifact.
@@
As a next step, we use this information to improve state preparation and measurement control. We calculated the new angular position of waveplates, considering the discovered retardance deviations, to implement the improved preparation and projection and repeated the experiment, obtaining $F_{\min}=0.991$, $\bar{F}=0.997$, $\bar{P}=0.995$, and $P_{\min}=0.983$. The results shown in Figure 2 indicate improved quality of state preparation and tomographic measurements. The small remaining artifacts are present due to the limited precision of the learned retardance deviations and also the repeatability of the waveplate rotation. Further, in Section S4 of the Supplementary material, we show the use of the presented method for classical rotating waveplate polarimetry [6].
@@
Conclusion - The mismatch between actual measurement operators and their theoretical counterparts assumed in reconstructing quantum states or devices introduces reconstruction artifacts. We showed how to leverage these artifacts to reveal the actual measurement operators and mitigate systematic errors in the tomography and state preparation. We experimentally applied this method to Pauli tomography of polarization-encoded photonic qubits and numerically analyzed several other measurement schemes. The proposed method makes the tomography independent of the measurement device. The main advantage is that the method does not require perfect knowledge or control of the probe states, which is virtually impossible to obtain. The method can calibrate already existing tomographic apparatus without the need for individual characterization of its constituent components. This is particularly useful in the case where the experimental setup is monolithic and individual components cannot be calibrated individually, e.g. in the case of integrated circuits. One can also calibrate the state preparation with the reversed version of the method. Furthermore, the presented approach applies locally on individual parties of a larger multiparty system; i.e. it is fully scalable. In summary, we developed the framework for accurate and scalable calibration of quantum measurement devices and also quantum state preparation, which is critical in any applications of quantum science and technology.
@@
The application of the reported measurement-device agnostic quantum measurement goes beyond the full quantum tomography. The same method will find its use for mitigation of measurement errors in various approximative and more scalable approaches such as Monte-Carlo sampling [36, 37], permutationally invariant tomography [38], compressed sensing [39, 40], etc. The method can also improve the accuracy of building and characterization of the photonic and quantum circuits in complex media [41, 42]. Finally, the reported approach also applies to experimental realizations of entanglement certification and quantification [23, 24, 22, 43, 25].
@@ Acknowledgements.
Acknowledgments - We acknowledge the support of the Czech Science Foundation under grant No. 21-18545S. M. B. acknowledges the support of Palacký University under grant No. IGA-PrF-2024-008. We acknowledge the use of cluster computing resources provided by the Department of Optics, Palacký University Olomouc. We thank J. Provazník for maintaining the cluster and providing support.
Supplemental Material: Measurement-device agnostic quantum tomography
I NOMAD setting in the numerical test
When we numerically investigated the case of general additive errors, we set the NOMAD [1] parameters as listed in Tab. 1. Please refer to the NOMAD 4 manual [2] for the meaning of the parameters. The optimized parameters were bounded to the interval (-29, 29) deg, i.e., roughly three standard deviations of the true parameter distribution. The seed was pseudo-randomly chosen at the time of optimization. The maximum size of the evaluation block was set as twice the problem dimension to allow optimal parallelization of the initial Latin hypercube search. We used the nomadlad [3] package to use NOMAD from Python. The necessary configuration was to set NOMAD to use multiple initial optimization points instead of just one by not specifying parameter X0 and setting the Latin hypercube search parameter (LH SEARCH) sufficiently high.
parameter value
BB OUTPUT TYPE OBJ
DIMENSION 12
BB MAX BLOCK SIZE 24
MAX BB EVAL 100 000
SEED random integer
DIRECTION TYPE ORTHO 2N
LH SEARCH 1100 1100
LOWER BOUND (-0.5, …, -0.5)
UPPER BOUND (0.5, …, 0.5)
Table 1: NOMAD (v4.1.0) settings used for the global optimization. Some parameters are not listed because they are not critical for this application.
II Artifact mitigation of multiplicative errors
Here, we explore the sampling strategies in the case of multiplicative errors. Since the artifact is state-dependent, the probe sampling strategy is important for minimizing the reconstruction artifacts.
As in the main text, all states are measured in projections onto eigenstates of Pauli operators. The projections are implemented by rotating the measured state using the following unitary operation $\hat{R}_{y}(\theta_{j})\hat{R}_{z}(\phi_{j})$ and subsequent projection of the resulting state onto state $|0\rangle$. The rotation angles $\theta_{j},\phi_{j}$ determine the selected projection and are listed in Tab 2.
We test a few sampling schemes and plot the cost function near the ground truth ($\delta^{\prime}=0.02$, $\epsilon^{\prime}=-0.04$) for each. The cost function is the peak-to-valley value of reconstructed probe state purities, $\Delta P=P_{\max}-P_{\min}$. The probe sets and cost functions are depicted in Figure S1. As probe states set, we selected eigenstates of Pauli operators (panels a,b), eight states in (panels c, d), icosahedron (panels e, f), and then a uniform sampling of Bloch sphere angular coordinates ($\theta,\phi$) with 14 states (panels g, h) and 22 states (panels i, j) and finally NASA HEALPIX [4, 5] with 108 probe states (panels k, l).
$j$ 1 2 3 4 5 6
$\theta_{j}$ ($y$-axis) 0 $\pi$ $\pi/2$ $\pi/2$ $\pi/2$ $\pi/2$
$\phi_{j}$ ($z$-axis) 0 0 $\pi$ 0 $\pi/2$ $23\pi/$ 2
Table 2: Rotation angles in radians for Pauli tomography. The analyzed state is first rotated by $\phi_{j}$ around the $z$-axis and $\theta_{j}$ around the $y$-axis and subsequently projected to state $|0\rangle$.
The sampling is sufficient when the cost function possesses a unique global minimum. We have run additional simulations, varied the ground truth, and observed the shape of the cost function. We tested each sampling scheme from Fig. S1. We sampled the cost function near the varying ground truth values of $\delta^{\prime}$ and $\epsilon^{\prime}$ and plotted its value as color maps, depicted in Fig. S1. The first suitable sampling scheme appears for eight probe states. Interestingly, for 14 probe states, there is a rift in the cost function landscape that causes ambiguity in the $\delta,\epsilon$ determination. The remaining sampling schemes exhibit a well-defined global minimum and therefore are suitable. The cost function landscapes in Fig. S1 are plotted for a single choice of $(\delta,\epsilon)$. To better characterize the sampling strategies, we repeated the same simulations for multiple ground truths. Figures S2-S6 show the simulated cost function for 8, 12, 17, 22, and 108 probe states. The ground truth is always plotted in the panel center. This additional characterization confirms the result – 8 probe states are enough and there is indeed anomalous behavior for 14 probe states. For 22 states, we sometimes see a local minimum, such as for ground truth parameters $\delta^{\prime}=0.15$, $\epsilon^{\prime}=0$. Due to this, the optimization has to be done carefully, ideally with several starting points, to find the global minimum of the cost function. In the case of 108 probe states, the cost function always had a single minimum.
The most interesting outcome of this analysis is that the success of the sampling strategy is not determined by the sheer number of probe states but rather by their suitable placement on the Bloch sphere. Generally, adding more probe states makes the error parameters determination more robust because we increase the obtained information, however, at the cost of longer measurement time. The exact cost function shape, sampling strategies, and shapes of the purity modulation depend on the imperfection type, its parameterization, and how we switch the measurement operators in the tomography.
Figure S1: Dependence of simulated artifact severity, represented as purity modulation amplitude depicted as 2D color maps in panels (a, c, …, k) on the selection of probe states set, depicted as points on projected Bloch sphere in panels (b, d, …, l). The figure shows the case for 6 (a,b), 8 (c,d), 12 (e,f), 14 (g,h), 22 (i,j), and 108 (k,l) probe states. The purity modulation amplitude is always plotted around the ground truth $\delta=0.02$, $\delta=-0.04$.
Figure S2: The panels show color-coded purity modulation as a function of assumed error parameters ($\delta,\epsilon$) observed in the set of eight reconstructed probe states. The purity modulation is sampled near the true value ($\delta^{\prime},\epsilon^{\prime}$) for each simulated ground truth. This true value is always plotted in the center of the panel. All panels share the color bar. To visually emphasize the presence of a single minimum, we use logarithmic normalization of color coding.
Figure S3: 12 probe states. The figure is arranged as Fig. S2.
Figure S4: 14 probe states. The figure is arranged as Fig. S2.
Figure S5: 22 probe states. The figure is arranged as Fig. S2.
Figure S6: 108 probe states. The figure is arranged as Fig. S2.
III Experimental setup details
We used a continuous-wave 810 nm fiber-coupled laser diode of 1 mW power. The laser diode was thermally stabilized using Peltier elements. The laser beam was decoupled into free space using an 11-mm aspheric lens and horizontally polarized using a Glan-Taylor prism. The angular positions of quarter-wave and half-wave plates controlled the resulting polarization of the beam. The following operator describes the action of a waveplate:
$\hat{W}_{\alpha,\gamma}=|\alpha\rangle\langle\alpha|+e^{-i\Gamma}|\alpha_{\bot}\rangle\langle\alpha_{\bot}|,$ (S1)
where $\alpha$ and $\Gamma$ are its angular position and retardance, and $|\alpha\rangle=\cos{\alpha}|0\rangle+\sin{\alpha}|1\rangle$ is a linearly polarized state, and $\alpha_{\bot}=\alpha+\pi/2$. To prepare a state located at colatitude $\theta$ and longitude $\phi$ on the Bloch sphere, we rotate the half- and quarter-wave plate to angles $x$ and $y$ given by the relations
$y=-\frac{1}{2}\arcsin\left(\sin\theta\sin\phi\right),$ (S2)
$x=\frac{1}{4}\arctan\left(\tan\theta\cos\phi\right)+\frac{1}{2}y+c(\theta),$ (S3)
Then we inserted a rotated quarter-wave plate that rotated the prepared state to demonstrate the robustness of our method to unitary operations in between the preparation and projection stages. The states were analyzed using half-waveplate and quarter-wave and subsequent projection onto horizontal and vertical polarization using a calcite beam-displacer. The waveplate settings for the Pauli tomography are listed in Tab. 3. All waveplates were mounted into a precision motorized rotation stage that provided bidirectional repeatability of ±$\pm 30$ mdeg.
The optical signals were measured by reading photo-currents from a pair of reverse-biased photodiodes. We normalized the readings using the sum of two signals to eliminate laser power fluctuation. The same results are obtained for a single-photon signal using two single-photon detectors.
The calibration step revealed the retardance deviation of the preparation and projection wave plates. To reconstruct the quantum states, we could directly use this knowledge to update our measurement operators. However, we take one extra step which allows us to prepare the states with greater precision and control the measurement operators precisely. We took these deviations into account in the verification step and calculated new angular positions of wave plates for tomographic projections and state preparation. Since Eq. S3 does not hold anymore, we used numerical optimization to find the desired wave plate settings.
To prepare a state $|\psi\rangle$, we find the wave plate angles $x,y$ using the following maximization:
$x,y=\mathrm{argmax}_{x,y}|\langle\psi|\hat{W}(x,\pi+\delta)\hat{W}(y,\pi/2+\epsilon)|0\rangle|^{2}.$ (S4)
$x,y=\mathrm{argmax}_{x,y}|\langle\langle 0|\hat{W}(x,\pi/2+\epsilon)\hat{W}(y,\pi+\delta)|\pi\rangle|^{2}.$ (S5)
state $|H\rangle$ $|V\rangle$ $\frac{|H\rangle+|V\rangle}{\sqrt{2}}$ $\frac{|H\rangle-|V\rangle}{\sqrt{2}}$ $\frac{|H\rangle+i|V\rangle}{\sqrt{2}}$ $\frac{|H\rangle-i|V\rangle}{\sqrt{2}}$
$j$ 1 2 3 4 5 6
$x_{j}$ (HWP) 0 $\pi/4$ $\pi/8$ $-\pi/8$ $-\pi/8$ $\pi/8$
$y_{j}$ (QWP) 0 0 0 0 $-\pi/4$ $\pi/4$
$x^{\prime}_{j}$ (HWP) 0 $\pi/4$ $\pi/8$ $-\pi/8$ $\pi/8$ $-\pi/8$
$y^{\prime}_{j}$ (QWP) 0 0 0 0 $\pi/4$ $-\pi/4$
Table 3: Waveplate angular position for Pauli tomography. The waveplates are placed as follows in the direction of the light propagation: quarter-wave plate and half-wave plate in the preparation, and half-wave plate and quarter-wave plate in the measurement stage. Primed symbols correspond to the state preparation, used in the reversed version of the method.
IV Rotating waveplate polarimetry
Consider the classical rotating quarter-wave plate polarimeter [6], depicted in Fig. S7(a), with the quarter-wave plate having the true retardance deviated by $\delta^{\prime}=5\deg$ from its nominal value. Moreover, consider pure probe states. The waveplate rotates continuously and effectively modulates the optical signal detected after the polarizer. Instead of likelihood maximization, we now use classical Fourier analysis to reconstruct the polarization state. We chose the Fourier analysis to show that our method is not limited to maximal likelihood reconstruction. We used eight probe states to detect the reconstruction artifact, and instead of purity modulation, we equivalently measured the severity with a modulation amplitude of the degree of polarization.
First, we treat the data with the ideal retardance ($\delta=0$) assumption. We see that the states from the ensemble suffer from artifacts. Here, we look at the intensity-normalized Stokes vector length $D=\frac{\sqrt{\sum\limits_{i=x,y,z}S_{i}}}{S_{0}}$, which is proportional to purity $P=\frac{1}{2}\left(1+D^{2}\right)$. Near state |$\ket{H}$, at the north pole of the Bloch sphere, the purity even exceeds 1 due to the reconstruction method. The modulation of $D$ is depicted in Fig. S7(c). This behavior is clearly a reconstruction artifact. Exceeding the unit purity is unsurprising because Fourier analysis does not guarantee physically sound results. We now vary $\delta$ assumed in reconstruction and observe the peak-to-peak value $\Delta D$ in the probe set. The dependence is depicted in Fig. S7(b). We see that the minimum lies at the true $\delta^{\prime}$. In panel (d), we show that the artifact disappears when the retardance deviation is taken into account.
Figure S7: Mitigation of artifact in rotating-waveplate polarimetry. (a) Measurement apparatus. A quarter-wave plate (purple) with retardance deviation $\delta^{\prime}=5$ deg is rotating in front of the linear polarizer. The photodiode signal is recorded and analyzed. (b) The minimal $\Delta D$ lies at the true retardance deviation $\delta^{\prime}$. (c) Initial mismatch results in state-dependent purity artifact. We plot the reconstructed probe states on the projected Bloch sphere and color-code their reconstructed stokes length $D=\sqrt{S^{2}_{x}+S^{2}_{y}+S^{2}_{z}}/S_{0}$. (d) When we considered the true retardance, the artifact vanished. Black circles depict the ideal testing states.
V Simulated experiment with path-encoded qubits on a photonic chip
As another example, we illustrate the measurement-device agnostic calibration for tomography of photonic on-chip path qubits [7, 8, 9, 10]. The qubit is encoded into the photon propagation in one or another waveguide (states $\ket{\uparrow}$ or $\ket{\downarrow}$), as depicted in Figure S8(a). The local heating of one waveguide controls the phase between the two paths. The heating is typically achieved by an electrical current flowing through a small heater element. A Mach-Zehnder interferometer serves as a variable coupler, and its interferometric phase controls the coupling ratio. Detection of photons in a single optical path implements a projective measurement of the path qubit. The circuit in Figure S8(a) implements a projection onto state |$\ket{\pi}$, whose colatitude and longitude on the Bloch sphere are controlled by optical phases. The relation between optical phases $\varphi_{1,2}$ and voltages $V_{1,2}$ applied on the heating elements is approximated with polynomials [8]
$\varphi_{1}(V_{1})=c_{1}+c_{2}V_{1}+c_{3}V_{1}^{2},$ (S6)
$\varphi_{2}(V_{2})=c_{4}+c_{5}V_{2}+c_{6}V_{2}^{2}.$ (S7)
Figure S8: Application of the method for on-chip photonic path qubit. (a) Projection apparatus for a path qubit. Applied voltage $V_{1,2}$ on heater elements (orange) causes thermal-based phase shift in the waveguide interferometers. These shifts determine the measurement projector. The photons in the upper output path are detected with a single-photon detector (SPAD). (b) 1D-section of 6D function $P_{\min}(c_{1},\dots,c_{6})$. The value of $P_{\min}$ at the initial assumption of $c_{j}$ is 0.955. Maximum lies at the true value $c^{\prime}_{6}$. (c) Initial mismatch results in state-dependent purity artifact. We plot probe states on the Bloch sphere and color-code their reconstructed purity. (d) When we consider the true retardance, the artifact vanishes. Parameters of the simulation are: $c^{\prime}_{1}=-1.887$, $c^{\prime}_{2}=0.105$ V-1, $c^{\prime}_{3}=0.05$ V-2, $c^{\prime}_{4}=-1.805$, $c^{\prime}_{5}=0.115$ V-1, and $c^{\prime}_{6}=0.05$ V-2. Initially assumed parameters were: $c_{1}=-1.750$, $c_{2}=0.100$ V-1, $c_{3}=0.050$ V-2, $c_{4}=-1.800$, $c_{5}=0.1105$ V-1, and $c_{6}=0.050$ V-2
Here, the voltages $V_{1,2}$ are controlled perfectly, but parameters $c_{j}$ are known only approximately. The other imperfections are neglected. We have increased the number of probe states to 22 to compensate for the higher number of optimization parameters. The simulated experiment, depicted in Fig. S8(c), shows the measurement operator mismatch artifact. We search the six-dimensional parametric space to find the point of minimal $\Delta P(c_{j})$ where the systematic error is mitigated. Panel (b) shows a one-dimensional section illustrating that the minimal $\Delta P$ corresponds to the true value $c^{\prime}_{6}$ (indicated by the dashed line). Taking the true values $c^{\prime}_{j}$ into account mitigates the reconstruction artifact, as shown in panel (d).

### v4

@@ VI Experiment: polarization state tomography
The action of the waveplate could be seen as the rotation of the Bloch sphere around axis $(\sin(2\alpha),0,\cos(2\alpha))$ with rotation angle $\Gamma$, where $\alpha$ determines the angular position of the waveplate and $\Gamma$ its retardance. The retardance usually slightly deviates from its nominal value. These deviations manifest as under- or over-rotation of the Bloch sphere. We parameterize the measurement operator by precisely controlled waveplate angular positions $x$, $y$, and half- and quarter-wave plate retardance deviations $\delta$ and $\epsilon$, respectively. These deviations are initially unknown to us, and we will use the presented method to estimate their values.
@@
We use eight probe states and perform the Pauli tomography for each. Initially, we assume zero deviations, $\delta=\epsilon=0$, and observe the purity modulation $\Delta P=0.0572(3)$. The number in parenthesis is one standard deviation at the last significant digit, e.g. $0.0572(3)=(572\pm 3)\cdot 10^{-4}$. The uncertainties were estimated using bootstrapping with the assumption of normally distributed photo-current with 0.1% relative standard deviation, which is the most dominant term and is related to the ammeter precision. The shot noise is negligible at the power level of our laser, compared to the technical noise. Then, we vary parameters $\delta$ and $\epsilon$ and update the measurement operators employed in the reconstruction to minimize the purity modulation. The optimal $\delta$ and $\epsilon$ are close to the true values because the reconstruction artifact is significantly reduced to $\Delta P=0.0140(2)$. We found the optimal parameters $\delta=4.5(2)$ deg, $\epsilon=-1.30(4)$ deg.
@@
We then reverse the process to determine retardance deviations $\tilde{\delta},\tilde{\epsilon}$ of the waveplates used in state preparation. We prepare the six eigenstates of Pauli operators, project each onto eight states corresponding to the original probe state set, and treat this data as before. We minimize the apparent purity modulation from 0.0700(3) to 0.0144(2) to find the optimal parameters $\tilde{\delta}=4.53(6)$ deg and $\tilde{\epsilon}=-3.58(3)$ deg.
The calibration step revealed the retardance deviation of the wave plates used for preparation and projection. To reconstruct the quantum states, we could directly use this knowledge to update our measurement operators. However, we take one extra step which allows us to prepare the states with greater precision and control the measurement operators precisely. We took these deviations into account in the verification step and calculated new angular positions of wave plates for tomographic projections and state preparation. Since Eq. 10 does not hold anymore, we used numerical optimization to find the desired wave plate settings ($x,y$):
$x,\,y=\mathrm{argmax}_{x,y}|\langle\psi|\hat{W}(x,\pi+\delta)\hat{W}(y,\pi/2+\epsilon)|0\rangle|^{2}.$ (11)
$x,\,y=\mathrm{argmax}_{x,y}|\langle\langle 0|\hat{W}(x,\pi/2+\epsilon)\hat{W}(y,\pi+\delta)|\pi\rangle|^{2}.$ (12)
At this point, we have information about actual measurement operators, and we have also calibrated the state preparation. This calibration was done in the presence of a unitary operation between state preparation and measurement, which was introduced by QWPX. This operation was effectively unknown because we did not use any prior information about QWPX during the process.
Now, we proceed to certify the correctness of such a calibration. The calibration should remain valid even if we change the central unitary operation introduced by QWPX. To test this, we removed the QWPX and then prepared an ensemble of 58 test states, depicted as empty circles in Figure 2(c,d).
As a reference, we first assumed perfect waveplates and performed state tomography for each test state. We indeed observed purity modulation $\Delta P=0.0607(2)$ with purity minimum $P_{\min}=0.9383(2)$. This modulation indicates reconstruction artifacts. The minimum fidelity of the reconstructed state to the target states was $F_{\min}=0.96688(4)$. These fidelity values indicate imprecise preparation.
In contrast, when we used our knowledge of $\delta$ and $\epsilon$ in the reconstruction and preparation, the minimum purity increased to $0.9830(3)$ and the minimum fidelity to $F_{\min}=0.9907(2)$. The information gained in the calibration clearly improved the reconstruction.
@@
The results in Figure 2 indicate improved quality of state preparation and tomographic measurements when we use the knowledge of $\delta$ and $\epsilon$ to improve both preparation and measurement. The small remaining artifacts are present due to the limited precision of the learned retardance deviations and also the repeatability of the waveplate rotation.
VII Path-encoded qubits on a photonic chip
Figure 3: Application of the method for on-chip photonic path qubit. (a) Projection apparatus for a path qubit. Applied voltage $V_{1,2}$ on heater elements (orange) causes thermal-based phase shift in the waveguide interferometers. These shifts determine the measurement projector. The photons in the upper output path are detected with a single-photon detector (SPAD). (b) 1D-section of 6D function $\Delta P(c_{1},\dots,c_{6})$. The value of $\Delta P$ at the initial assumption of $c_{j}$ is $1.8\cdot 10^{-2}$ on the ensemble of probe states. Minimum $\Delta P=4.5\cdot 10^{-4}$ lies at the true value $c^{\prime}_{6}$. (c) Initial mismatch results in state-dependent purity artifact. We plot test states on the Bloch sphere as circles and color-code their reconstructed purity. (d) When we consider the true retardance, the artifact vanishes.
As another example, we numerically illustrate the measurement-device agnostic calibration for tomography of photonic on-chip path qubits [30, 35, 36, 37]. The qubit is encoded into the photon propagation in one or another waveguide (states $\ket{\uparrow}$ or $\ket{\downarrow}$), as depicted in Figure 3(a). The local heating of one waveguide controls the phase between the two paths. The heating is typically achieved by an electrical current flowing through a small heater element. A Mach-Zehnder interferometer serves as a variable coupler, and its interferometric phase controls the coupling ratio. Detection of photons in a single optical path implements a projective measurement of the path qubit. The circuit in Figure 3(a) implements a projection onto state |$\ket{\pi}$, whose colatitude and longitude on the Bloch sphere are controlled by optical phases. The relation between optical phases $\varphi_{1,2}$ and voltages $V_{1,2}$ applied on the heating elements is approximated with polynomials [35]
$\varphi_{1}(V_{1})=c_{1}+c_{2}V_{1}+c_{3}V_{1}^{2},$ (13)
$\varphi_{2}(V_{2})=c_{4}+c_{5}V_{2}+c_{6}V_{2}^{2}.$ (14)
Here, the voltages $V_{1,2}$ are controlled perfectly, but parameters $c_{j}$ are known only approximately. The other imperfections are neglected. We have increased the number of probe states to 22 to compensate for the higher number of optimization parameters. The simulated experiment, depicted in Fig. 3(c), shows the measurement operator mismatch artifact. We search the six-dimensional parametric space to find the point of minimal $\Delta P(c_{j})$ where the systematic error is mitigated. Panel (b) shows a one-dimensional section illustrating that the minimal $\Delta P$ corresponds to the true value $c^{\prime}_{6}$. Taking the true values $c^{\prime}_{j}$ into account mitigates the reconstruction artifact, as shown in panel (d).
VIII Conclusion
@@
The mismatch between actual measurement operators and their theoretical counterparts assumed in reconstructing quantum states or devices introduces reconstruction artifacts. We showed how to leverage these artifacts to reveal the actual measurement operators and mitigate systematic errors in the tomography and state preparation. We experimentally applied this method to Pauli tomography of polarization-encoded photonic qubits and numerically analyzed several other measurement schemes. The proposed method makes the tomography independent of the measurement device. The main advantage is that the method does not require perfect knowledge or control of the probe states, which is virtually impossible to obtain. The method can calibrate already existing tomographic apparatus without the need for individual characterization of its constituent components. This is particularly useful in the case where the experimental setup is monolithic and individual components cannot be calibrated individually, e.g. in the case of integrated circuits. One can also calibrate the state preparation using the same method. Furthermore, the presented approach applies locally to individual parties of a larger multiparty system; i.e. it is fully scalable.
In summary, we developed the framework for accurate and scalable calibration of quantum measurement devices and also quantum state preparation, which is critical in any applications of quantum science and technology.
@@
The application of the reported measurement-device agnostic quantum measurement goes beyond the full quantum tomography [13, 22]. The same method will find its use for mitigation of measurement errors in various approximative and more scalable approaches such as Monte-Carlo sampling [38, 39, 40, 41], permutationally invariant tomography [42], compressed sensing [43, 44], etc. The method can also improve the accuracy of building and characterization of the photonic and quantum circuits in complex media [45, 46]. Finally, the reported approach also applies to experimental realizations of entanglement certification and quantification [26, 27, 25, 47, 28].
@@ Acknowledgements.
We acknowledge the support of the Czech Science Foundation under grant No. 21-18545S. M. B. acknowledges the support of Palacký University under grant No. IGA-PrF-2024-008. We acknowledge the use of cluster computing resources provided by the Department of Optics, Palacký University Olomouc. We thank J. Provazník for maintaining the cluster and providing support.